# Kiet A. NguyenAdheesh JuvekarTianjiao YuMuntasir WahedIsmini Lourentzou

### v1

@@ 6 Conclusion
This paper introduces the novel task of part-focused semantic co-segmentation, which involves the segmentation of common and unique parts across multiple images, laying the groundwork for future research in enhancing the capability of Large Vision-Language Models (LVLMs) to analyze and interpret complex visual data in a granular manner. To solve this task, we propose Calico, an LVLM that incorporates a novel correspondence extraction module and an adaptation module to handle multi-image part relationships. Experiments conducted on the newly curated MixedParts dataset, demonstrate that Calico can effectively identify and segment common/unique parts with high accuracy, outperforming existing models.
@@ Acknowledgments
This research is based upon work partially supported by U.S. DARPA ECOLE Program No. HR00112390062. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.1 Single-Image Part Segmentation Datasets
    PartImageNet [15] is a high-quality part-focused extension of ImageNet [10] covering a variety of object classes with mostly animals (e.g. bird, fish, etc.) to facilitate non-rigid part understanding. Each image in PartImageNet contains only one foreground object, which can encourage the model to focus on important foreground objects when comparing two images. Prior to use, we make modifications to the original dataset to enhance generalizability, as detailed in Appendix D.
@@
    ADE20K-Part234 [61] is a revised version of the ADE20K scene parsing dataset [73, 74] with an emphasis on object parts. The original ADE20K dataset encompasses a wide variety of scenes, including indoor, outdoor, and urban environments, which naturally lends itself to more complex visual signals covering multiple objects in contrast to PartImageNet. However, less than $15\%$ of the dataset contains part annotations. In addition, some part labels are too granular, which may encourage overfitting while not being too beneficial for general part understanding (e.g. “table stretcher” and “table h-stretcher”). To amend these disadvantages, Wei et al. [61] introduces ADE20K-Part234, a clean, part-focused version of ADE20K for improved part analysis.
@@
    PACO-LVIS [45], likewise, is a component-centric version of LVIS [14] that is based on COCO [30] and focuses on diverse common everyday objects, further contributing to the diversity of object categories in MixedParts. PACO contains an extensive list of object-part categories as well as complex images with multiple objects and parts, providing finer granularity for part understanding.
A.2 Constructing MixedParts
To build MixedParts, we first select 1,885 pairs of object categories across all 3 datasets that have at least one common part label (e.g. “armchair’s seat” and “swivel chair’s seat”) to ensure annotation availability. However, due to the flexibility of natural language, the same words may be used to describe parts of objects that are not typically intuitively comparable. For example, even though both a bus and a microwave oven may have a door as a constituent part, they are not commonly compared. Therefore, we manually curate intuitively comparable object pairs from all possible pairs, resulting in 964 pairs of categories across all 3 datasets.
@@ A.2 Constructing MixedParts
Figure 6: MixedParts Dataset Overview.
Statistic Task Common Objects Common Parts Unique Parts
Total # Pairs 1,257,247 1,724,937 5,568,099
Average # Pairs/Sample 1.049 2.019 5.272
Maximum # Pairs/Sample 5 28 72
Median # Pairs/Sample 1 2 3
Table 4: MixedParts statistics on # pairs of objects/parts. Each sample in MixedParts corresponds to an image pair. For unique parts, pair corresponds to one unique part present in a single image instead of both images, thus explaining the higher number of instances.
@@ Appendix B MixedParts Dataset Statistics
We provide details for the datasets used in constructing MixedParts. Figure 6 and Table 4 offer a comprehensive overview of the dataset, illustrating its structure and the distribution of object and part pairings. The inner circle of the donut chart represents the entire dataset, comprising a total of 2,386,746 samples (image pairs), segmented into three primary sources: PartImageNet, PACO-LVIS, and ADE20K-Part234, each contributing roughly a third to the total dataset. The outer ring of the chart further divides these sources into categories of objects and parts, with each segment roughly occupying half of its respective section, ensuring a balanced representation of various part-object relationships. In MixedParts, there are 1,257,247 pairs of common objects, 1,724,937 pairs of common parts, and 5,568,099 unique parts. On average, each sample comprises approximately 1 common object, 2 common parts, and 5.3 unique parts. The maximum number of pairs per sample is 5 for common objects, 28 for common parts, and 72 for unique parts, indicating a wide range in the complexity of the sample pairings. The median number of pairs per sample is 1 for common objects, 2 for common parts, and 3 for unique parts, reflecting that the majority of samples contain a single pair of common objects and more than 1 common/unique part, since one object pair may consist more than 1 common and unique parts. MixedParts consists of 141 object categories and 199 part categories covering a wide variety of objects and parts with 69,911 object instances and 159,140 part instances. Figures 8-9 present the top-30 object and part distributions in the dataset, while Figure 10 and Figure 11 present the distributions of top-30 common and unique parts, respectively.
@@ C.1 Implementation Details
We use PyTorch [42] to implement and optimize Calico and DeepSpeed [47] for efficient training. We initialize our model on GLaMM’s [46] FullScope checkpoint. [^3] https://huggingface.co/MBZUAI/GLaMM-FullScope. All training is conducted on four NVIDIA A40 GPUs with 48GB memory. LoRA layers were set with a rank of 8 and a scaling factor of 16. We train for 10 epochs, each comprising 500 steps, using the AdamW optimizer [36] with 3e-4 initial learning rate and beta coefficients set to 0.9 and 0.95, respectively. We employ a dropout rate of 0.05, 1.0 gradient clipping, and set the batch size to 4. The loss coefficients are set to $\lambda_{\text{text}}=1.0$, $\lambda_{\text{focal}}=2.0$, and $\lambda_{\text{Dice}}=0.5$.
Baselines. We provide implementation details for our zero-shot VLPart baseline [56], as well as the fine-tuned LISA [24] and GLaMM [46] baselines. For VLPart, we use the default parameters from the official repository [^4] https://github.com/facebookresearch/VLPart., including the detection confidence threshold of 0.7 (i.e., retaining predictions with confidence scores above 0.7) to initialize the model. For a given two-image data sample, we perform zero-shot inference on both images and analyze all predicted object and part masks. For common objects, we compare the sets of predicted objects across the images. For common and unique parts, we aggregate the sets of all predicted parts and derive the associated objects from these parts. For GLaMM and LISA, which are not natively designed to handle multiple images, we adapt the code to distinguish segmentation tokens belonging to different images. This is achieved by appending image identifiers to the [SEG] tokens, e.g., (IMAGE1), enabling the models to process the image-specific token sets separately.
@@ C.2 Evaluation Metrics
Our evaluation consists of two parts: 1) evaluating the segmentation performance and 2) assessing the performance in generating semantic labels for objects and parts. To evaluate the performance of models on segmentation tasks, we employ mean Intersection over Union (mIoU), a widely used metric that effectively measures how much the predicted segmentation masks overlap with the ground truth masks across a dataset. We also employ Average Precision at a 50% IoU threshold (AP50), a metric for evaluating segmentation masks that measures the average precision across all recall levels at a fixed IoU threshold. AP50 considers a predicted mask correct if its IoU with the ground truth mask is at least 50%, summarizing model performance at this specific threshold. Additionally, following recent work [46], we leverage a mask Recall metric, which evaluates region-specific grounding by using a two-phase validation approach based on IoU and SentenceBERT [48] similarity thresholds. Specifically, given a ground truth match, we choose the prediction with the highest IoU surpassing a threshold of 50% that also has a text similarity score higher than 50% as a True Positive for recall computation.
Task Method mIoU AP50 Recall SS S-IoU
Common Objects Multi-Image VLPart [56] 15.8 41.3 44.3 57.8 53.7
Multi-Image GLaMM [46] 49.2 61.6 59.4 73.6 72.0
Multi-Image LISA [24] 49.6 62.8 60.4 78.2 76.9
Calico (ours) 58.3 68.8 68.9 88.5 87.6
Common Parts Multi-Image VLPart [56] 15.5 43.4 28.4 53.9 38.8
Multi-Image GLaMM [46] 24.8 45.5 33.7 58.1 46.4
Multi-Image LISA [24] 26.2 46.8 37.3 61.2 51.9
Calico (ours) 30.3 50.1 42.9 68.4 57.9
Unique Parts Multi-Image VLPart [56] 6.0 41.4 29.4 64.1 43.7
Multi-Image GLaMM [46] 16.8 47.6 29.4 60.0 50.5
Multi-Image LISA [24] 16.3 46.8 29.1 59.6 50.9
Calico (ours) 22.8 52.7 39.7 72.7 65.1
MixedParts Multi-Image VLPart [56] 12.4 42.0 34.0 58.6 45.4
Multi-Image GLaMM [46] 30.3 51.6 40.8 63.9 56.3
Multi-Image LISA [24] 30.7 52.1 42.3 66.3 59.9
Calico (ours) 37.1 57.2 50.5 76.5 70.2

### v2

@@ 6 Conclusion
This paper introduces the novel task of part-focused semantic co-segmentation, which involves the segmentation of common and unique parts across multiple images, laying the groundwork for future research in enhancing the capability of Large Vision-Language Models (LVLMs) to analyze and interpret complex visual data in a granular manner. To solve this task, we propose Calico, an LVLM that incorporates a novel correspondence extraction module and an adaptation module to handle multi-image part relationships. Experiments conducted on the newly curated MixedParts dataset demonstrate that Calico can effectively identify and segment common/unique parts with high accuracy, outperforming existing models.
@@ Acknowledgments
This research is based upon work supported by U.S. DARPA ECOLE Program No. HR00112390062 and U.S. DARPA SciFy Program No. HR001125C0303. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein. This work used the Delta GPUs provided by the National Center for Supercomputing Applications through allocations CIS240752, CIS240817, and CIS250130 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS, Boerner et al. [4]) program, supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@ A.1 Single-Image Part Segmentation Datasets
    PartImageNet [18] is a high-quality part-focused extension of ImageNet [13] covering a variety of object classes with mostly animals (e.g. bird, fish, etc.) to facilitate non-rigid part understanding. Each image in PartImageNet contains only one foreground object, which can encourage the model to focus on important foreground objects when comparing two images. Prior to use, we make modifications to the original dataset to enhance generalizability, as detailed in Appendix A.2.
@@
    ADE20K-Part234 [65] is a revised version of the ADE20K scene parsing dataset [77, 78] with an emphasis on object parts. The original ADE20K dataset encompasses a wide variety of scenes, including indoor, outdoor, and urban environments, which naturally lends itself to more complex visual signals covering multiple objects in contrast to PartImageNet. However, less than $15\%$ of the dataset contains part annotations; in addition, some part labels are too granular, which may encourage overfitting while not being too beneficial for general part understanding (e.g. “table stretcher” and “table h-stretcher”). To amend these disadvantages, Wei et al. [65] introduces ADE20K-Part234, a clean, part-focused version of ADE20K for improved part analysis.
@@
    PACO-LVIS [49] is a part-centric version of LVIS [17] that is based on COCO [34] and focuses on diverse everyday objects, further contributing to the diversity of object categories in MixedParts. PACO contains an extensive list of object-part categories as well as complex images with multiple objects and parts, providing finer granularity for part understanding.
To curate MixedParts, we first select 1,885 pairs of object categories across all 3 datasets that have at least one common part label (e.g. “armchair’s seat” and “swivel chair’s seat”) to ensure annotation availability. However, due to the ambiguity of natural language, the same part name can refer to object parts that are not typically intuitively comparable. For example, even though both a bus and a microwave oven may have a door, they are not commonly compared. Therefore, we manually curate intuitively comparable object pairs from all possible pairs, resulting in 964 pairs of categories across all 3 datasets.
@@
A.2 PartImageNet Modifications
While PartImageNet provides high-quality segmentation masks across a diverse range of object classes, its categories are often too abstract and not commonly used in natural language due to their broad scope. For instance, PartImageNet classifies all four-legged animals under the “quadruped” supercategory. Although technically accurate, such classifications are too generic for practical use in everyday language.
To address this, we manually selected the most commonly used object class associated with each category provided by ImageNet. The dataset includes WordNet synset IDs (e.g., “n02071294”) that correspond to specific object classes. Using these synset IDs, we extracted the hierarchical path of each class within the WordNet graph. For example, the synset ID “n02071294” maps to the following WordNet path: living_thing $\to$ organism $\to$ animal $\to$ chordate $\to$ vertebrate $\to$ mammal $\to$ placental $\to$ aquatic_mammal $\to$ cetacean $\to$ whale $\to$ toothed_whale $\to$ dolphin $\to$ killer_whale. From this path, we selected whale as the representative object category for this class. We repeat this process for all object classes provided by PartImageNet. The object classes we use alongside the original supercategory in parentheses are shown in Table 6.
@@ Appendix B MixedParts Dataset Statistics
Figure 9 and Table 3 provide a comprehensive overview of our benchmark, illustrating its structure and the distribution of objects and parts. The inner circle of the donut chart represents the entire dataset, comprising a total of 2,382,747 samples (image pairs), sourced from PartImageNet, PACO-LVIS, and ADE20K-Part234, each contributing to roughly a third of the total dataset. The outer ring of the chart further divides these sources into categories of objects and parts, contributing roughly half, ensuring balanced representation of various object–part relationships. In MixedParts, there are 2,509,552 instances of common objects, 3,443,758 instances of common parts, and 5,557,188 instances of unique parts. On average, each image pair includes 2 object instances (one per image), 4 common part instances, and 5.3 unique part instances. Sample complexity varies widely, with a maximum of 10 common objects, 56 common parts, and 72 unique parts per sample. Median counts are 2 for common objects, 4 for common parts, and 3 for unique parts, reflecting that a single object pair can involve multiple shared and unique parts. The dataset spans 141 object categories and 196 part categories, covering a wide variety of objects and parts, totaling 65,960 distinct object instances and 154,463 distinct part instances. Figures 11-12 present the top-30 most frequent object and part categories, while Figures 13-14 show the top-30 common and unique parts, respectively.
@@ C.1 Implementation Details
We use PyTorch [46] to implement and optimize Calico, with DeepSpeed [51] for efficient training. We initialize our model on GLaMM’s [50] pretrained checkpoint, [^1] https://huggingface.co/MBZUAI/GLaMM-GranD-Pretrained. the Q-Former module on InstructBLIP’s [10] Vicuna-7B-aligned checkpoint, [^2] https://storage.googleapis.com/sfr-vision-language-research/LAVIS/models/InstructBLIP/instruct_blip_vicuna7b_trimmed.pth. and align them with 3.2M samples from GLaMM’s GranD dataset. All training is conducted on four NVIDIA A40 GPUs with 48GB memory. LoRA layers are applied to the query and value projections, configured with a rank of 8 and a scaling factor of 16. We train for 10 epochs, each comprising 500 steps, using the AdamW optimizer [40] with a 4e-4 initial learning rate and beta coefficients set to 0.9 and 0.95, respectively. We warm up the learning rate over 100 training steps and then use a linear decay scheduler over the remaining training time. We employ a dropout rate of 0.05, 1.0 gradient clipping, and set the batch size to 4 alongside gradient accumulation every 4 steps. The loss coefficients are set to $\lambda_{\text{text}}\!=\!1.0$, $\lambda_{\text{focal}}\!=\!2.0$, and $\lambda_{\text{Dice}}\!=\!0.5$. We retain all original implementation details for our baselines while keeping the effective batch size the same across all finetuned approaches to ensure fairness. We use DeepSpeed’s profiler [^3] https://www.deepspeed.ai/tutorials/flops-profiler/. to compute TFLOPS.
Figure 9: MixedParts Dataset Overview.
Statistic Common Objects Common Parts Unique Parts
Total # Instances 2,509,552 3,443,758 5,557,188
Average # Instances/Sample 2.099 4.037 5.269
Maximum # Instances/Sample 10 56 72
Median # Instances/Sample 2 4 3
Table 3: MixedParts Statistics on # instance of objects/parts. Here, an object/part instance in MixedParts refers to a single occurrence of an object or part in a sample, while a sample corresponds to an image pair.
Baselines. We provide implementation details for our baselines, which include three zero-shot and two finetuned LVLM-based approaches:
[…]
    Cascade: To evaluate whether a simple modular pipeline can effectively solve part-focused semantic co-segmentation, we first design an cascaded approach that sequentially integrates strong zero-shot models, each excelling at different subcomponents of the task: (1) Sparkles [22], a powerful multi-image model, is used to identify relevant objects across images, followed by (2) GPT-4o API [1] to extract object-part relationships, and finally, (3) LISA [27], a segmentation-based LVLM, to produce segmentation masks for the identified objects and parts. Given an image pair, we first prompt Sparkles to list the objects or parts present across both images. Next, we prompt GPT-4o with this information to identify the common or unique objects or parts, depending on the task. Finally, we prompt LISA with the objects/parts extracted by GPT-4o to generate pixel-level segmentation masks. Prompt templates for all models of this cascaded approach are shown in Figure 10.
[…]
    Multi-Image VLPart: VLPart [60] is an open-vocabulary part segmentation model that can segment object parts at different granularities. It combines a conventional Mask R-CNN [19] with a modern Swin Transformer backbone [38] and a CLIP [48] text classifier for open-world image classification. Although VLPart cannot perform co-segmentation, its strong zero-shot object-part segmentation capabilities enable a simple baseline: performing segmentation on individual images and simply examining the common and unique predictions. Specifically, for a given image pair, we perform inference on both images and extract all predicted object and part masks. To identify common objects, we compare the sets of predicted objects across images. For common and unique parts, we aggregate the sets of all predicted parts and derive their associated objects. We follow the default configuration from the official repository, [^4] https://github.com/facebookresearch/VLPart. including the detection confidence threshold of 0.7 (i.e., retaining predictions with confidence scores above 0.7) to initialize the model. We use the checkpoint with a Swin-base cascade backbone trained on the parsed ImageNet dataset. [^5] https://github.com/PeizeSun/VLPart/releases/download/v0.1/swinbase_cascade_lvis_paco_pascalpart_partimagenet_inparsed.pth.
[…]
    Multi-Image PartGLEE: PartGLEE [32] is a recent foundation model that uses Q-Former to query object and part representations from large-scale training data, boasting better object-part segmentation capabilities than VLPart. For this baseline, we simply obtain single-image outputs from PartGLEE (100 masks per image), keeping only those with non-zero scores (55 on average). Similar to Multi-Image VLPart, we compare the predicted class labels across the image pair to identify common and unique objects and parts. We use PartGLEE’s official implementation, [^6] https://github.com/ProvenceStar/PartGLEE.git. initializing with the released Swin-Large checkpoint and default settings.
[…]
    Multi-Image LISA: LISA [27] is an LVLM trained to perform referring segmentation, built on a Vicuna-7B backbone and SAM. Since LISA was not trained for multi-image processing, we replicate Calico’s multi-image implementation on the LISA codebase. We initialize LISA on their 7B checkpoint [^7] https://huggingface.co/xinlai/LISA-7B-v1. and finetune the mask decoder alongside LoRA.
[…]
    Multi-Image GLaMM: GLaMM [50] is a strong single-image segmentation-based LVLM constructed for the Grounded Conversation Generation (GCG) task involving multi-turn pixel-grounded dialogue. GLaMM combines a Vicuna-based language backbone with SAM, similarly to LISA, and a novel RoIAlign-based region encoder. We implement multi-image processing for GLaMM by sequentially feeding the mask decoder with the encoded images and corresponding segmentation tokens to obtain segmentation masks. Both Calico and Multi-Image GLaMM are initialized from GLaMM’s pretrained model weights. [^8] https://huggingface.co/MBZUAI/GLaMM-GranD-Pretrained. Since LISA and GLaMM are not natively designed to handle multiple images, we adapt Calico’s implementation for distinguishing segmentation tokens belonging to different images. This is achieved by appending image identifiers to the [SEG] tokens, e.g., (IMAGE1), enabling the models to process the image-specific token sets separately.
@@ C.2 Evaluation Metrics
Our evaluation consists of two parts: (1) segmentation quality and (2) semantic label accuracy for objects and parts. To evaluate performance on segmentation tasks, we employ mean Intersection over Union (mIoU), a widely used metric that measures the average overlap between predicted and ground truth masks across all classes. We also employ Average Precision at a 50% IoU threshold (AP50), which measures the average precision across all recall levels at a fixed IoU threshold. Specifically, AP50 considers a predicted mask correct if its IoU with the ground truth mask is at least 50%, summarizing model performance at this specific threshold. Additionally, following GLaMM [50], we report Recall, which evaluates region-specific grounding by using a two-phase validation approach based on IoU and SentenceBERT [52] similarity thresholds. In particular, a prediction is considered a true positive if its mask has IoU $\geq$ 50% with the ground truth and its text label exceeds 50% similarity with the ground truth.
To evaluate semantic labeling performance for the segmented objects and parts, we compute Semantic Similarity and Semantic IoU. Semantic Similarity (SS) measures similarity between predicted and true labels in the SentenceBERT [52] embedding space, following prior work [71, 9], while Semantic IoU (S-IoU) computes token-level overlap between predicted and ground truth labels, defined as
[…]
where $V(y)$ denotes the set of words comprising label $y$.
[…]
CO CP UP All
Multi-Image GLaMM [50] 54.9 576 42.3 8.4 21.3 26.3 18.7
Multi-Image LISA [27] 55.5 256 44.0 5.5 14.9 18.9 13.1
Calico 59.7 32 28.5 4.3 10.2 12.3 9.1
Table 4: Calico Efficiency in Numbers. Alongside TFLOPS, we show inference time per sample in seconds for the common objects (CO), common parts (CP), unique parts (UP) tasks, as well as the average (All).
Appendix D Additional Experiments
D.1 Computational Efficiency Evaluation
For multi-image tasks, computational efficiency is critical to ensure scalability in inference and training as the number of images increases. To address this, Calico employs Q-Former to query image embeddings, reducing the number of tokens per image to just 32, which is 8 times fewer tokens than LISA (256 tokens) and 18 times fewer than GLaMM (576 tokens). As shown in Table 4, this reduces TFLOPS by 35.23% over LISA and 32.62% over GLaMM (LISA:44.0 vs. GLaMM:42.3 vs. Calico:28.5 TFLOPS), yielding a $\sim$1.5$\times$×speed-up and a 7.6% relative performance gain over LISA. These results demonstrate that Calico offers substantial improvements in both computational efficiency and task performance compared to strong baselines.
D.2 Per-Task Experimental Results
Table 5 presents results decomposed into the 3 MixedParts subtasks (common objects, common parts, and unique parts). The decreasing performance across all models delineates the incremental difficulty of the tasks, i.e., with common objects being the easiest task, followed by common parts, and finally unique parts as the most challenging task. Calico consistently outperforms all baselines across all 3 tasks, with strong improvements in scores across all metrics compared to the next best baseline.
Task Method AP50 mIoU Recall SS S-IoU
Common Objects Cascade [22, 1, 27] 7.9 37.0 28.2 38.1 29.6
Multi-Image PartGLEE [32] 1.5 33.5 9.7 87.2 82.8
Multi-Image VLPart [60] 18.2 42.4 45.6 58.8 55.2
Multi-Image GLaMM [50] 63.2 71.6 73.3 86.5 86.2
Multi-Image LISA [27] 60.2 70.0 71.9 86.0 85.3
Calico 69.2 75.2 78.5 93.7 93.4
Common Parts Cascade [22, 1, 27] 3.5 18.6 11.7 25.6 6.5
Multi-Image PartGLEE [32] 1.9 32.0 10.5 78.4 63.3
Multi-Image VLPart [60] 15.7 44.5 29.1 54.4 39.9
Multi-Image GLaMM [50] 36.7 52.7 48.4 68.5 59.0
Multi-Image LISA [27] 37.3 54.2 50.5 73.9 63.4
Calico 38.4 56.9 51.9 74.4 64.4
Unique Parts Cascade [22, 1, 27] 5.7 28.1 17.2 32.8 8.2
Multi-Image PartGLEE [32] 0.1 22.3 8.8 69.9 43.9
Multi-Image VLPart [60] 6.4 41.5 29.0 64.0 44.4
Multi-Image GLaMM [50] 28.9 55.3 43.1 75.3 68.4
Multi-Image LISA [27] 26.6 55.0 44.1 76.2 68.7
Calico 30.2 59.1 48.8 80.1 73.4
MixedParts Cascade [22, 1, 27] 5.7 27.9 19.0 32.2 14.8
Multi-Image PartGLEE [32] 1.2 29.3 9.7 78.5 63.3
Multi-Image VLPart [60] 13.4 42.8 34.6 59.1 46.5
Multi-Image GLaMM [50] 42.9 59.9 54.9 76.8 71.2
Multi-Image LISA [27] 41.4 59.7 55.5 78.7 72.5
Calico 45.9 63.7 59.7 82.7 77.1